# Probing Hidden Topology with Quantum Detectors

### v1

@@ I Introduction
Black holes are amongst the most fascinating objects in nature, in large part because they cloak certain regions of space and time from all distant observers. While black holes created in a gravitational collapse are not time symmetric, their eternal counterparts contain both a future horizon and past horizon, and continuing the spacetime analytically past these horizons reveals new structure, a prime example of which is the second asymptotically flat region in the Kruskal(-Szekeres-Fronsdal) [1, 2, 3] extension of positive mass Schwarzschild spacetime. The structure past the horizon has a pivotal role in recognising Hawking temperature as the distinguished temperature for quantum fields in thermal equilibrium with the black hole [4], and it provides the cornerstone to the view of black hole entropy as a fundamentally quantum-gravitational concept [5].
@@ III.4 Swedish geon
The Swedish geon [17] is a single-exterior black-and-white hole quotient of $\text{AdS}_{3}$, with the exterior metric given by the BTZ exterior (14), but with behind-the-horizon identifications that give the spacetime the spatial topology of a punctured torus, the puncture being at the AdS infinity, $r\to\infty$ in (14). We review the construction of this spacetime in the three appendices, showing that the spacetime is a quotient of a subset of $\text{AdS}_{3}$ by the free group generated by the $O(2,2)$ matrices $A$ and $B$ given by
@@
IV Response Rate Comparisons
@@
In this section we present the image sum expressions for the detector’s transition rate on the BTZ black hole, the $\mathbb{R}\text{P}^{2}$ geon and the Swedish geon.
@@
IV.1 Image sum in the state induced from the $\text{AdS}_{3}$ global vacuum
@@
On each of the three quotients of $\text{AdS}_{3}$, we consider the state induced from the $\text{AdS}_{3}$ global vacuum. Denoting by $G$ the discrete group by which the quotient is taken, the Wightman function is given by the method-of-images sum
@@
In all three spacetimes, we consider a UDW detector on the static exterior wordline that is given in the BTZ exterior coordinates (14) by
@@ IV.2 BTZ
IV.3 $\mathbb{R}\text{P}^{2}$ geon
@@ Acknowledgements.
We thank Fredrik Strömberg for discussions on fundamental domains. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.
@@ Appendix C Topology
As the fundamental group of the punctured torus is the free group generated by the two identifications shown in Figure 8(a) [33, Section 1.2][34], it follows that $A/\{\pm 1\}$ and $B/\{\pm 1\}$ provide a pair of free generators for the action of the fundamental group on the Poincaré disc.

### v2

@@ I Introduction
Black holes are amongst the most fascinating objects in nature, in large part because they cloak certain regions of space and time from all distant observers. While black holes created in a gravitational collapse are not time symmetric, their eternal counterparts contain both a future horizon and past horizon, and continuing the spacetime analytically past these horizons reveals new structure, a prime example of which is the second asymptotically flat region in the Kruskal(-Szekeres-Fronsdal) [1, 2, 3] extension of positive mass Schwarzschild spacetime. The structure past the horizon has a pivotal role in recognising Hawking temperature as the distinguished temperature for quantum fields in thermal equilibrium with the black hole [4, 5], and it provides the cornerstone to the view of black hole entropy as a fundamentally quantum-gravitational concept [6].
@@ III.4 Swedish geon
The Swedish geon [18] is a single-exterior black-and-white hole quotient of an open region in $\text{AdS}_{3}$, with the exterior metric given by the BTZ exterior (14), but with behind-the-horizon identifications that give the spacetime the spatial topology of a punctured torus, the puncture being at the AdS infinity, $r\to\infty$ in (14). We review the construction of this spacetime in the three appendices, showing that the spacetime is a quotient of a subset of $\text{AdS}_{3}$ by the free group generated by the $O(2,2)$ matrices $A$ and $B$ given by
@@
IV Transition rate comparisons
@@
In this section we present the image sum expressions for the detector’s transition rate on the BTZ black hole, the $\mathbb{R}\text{P}^{2}$ geon and the Swedish geon, in the state induced from the $\text{AdS}_{3}$ global vacuum.
IV.1 Preamble: choice of the state
To begin, recall that Fock quantisation of a real Klein-Gordon field relies on a decomposition of the complexified classical solution space into a positive Klein-Gordon norm subspace and a negative Klein-Gordon norm subspace. In Minkowski spacetime, choosing the positive norm subspace to be the positive frequency subspace with respect to Minkowski time translations results in the standard Minkowski Fock vacuum as the ground state of the theory, and this vacuum is physically interpreted as the no-particle state for all inertial observers [30].
In curved spacetime, by contrast, there is no general criterion for choosing the positive norm subspace. In curved spacetimes with symmetries or asymptotic symmetries, however, the symmetries may offer a criterion for this choice, reflecting a choice of physical input. An example is de Sitter space, where de Sitter invariance combined with technical regularity properties singles out the Fock vacuum known as the Euclidean, Chernikov-Tagirov or Bunch-Davies vacuum [31, 32, 33], which has a physical interpretation as a thermal state for inertial observers [34]. Another example is an expanding cosmology with an early time asymptotically de Sitter era, where the asymptotic early time symmetries single out a Bunch-Davies state [35].
An example closer to the spacetimes considered in the present paper is the Schwarzschild black hole spacetime. In the exterior region, Schwarzschild time translations single out the state known as the Boulware state [36], which is close to the Minkowski vacuum at asymptotically flat infinity, but is singular on approaching the black and white hole horizons. On the Kruskal extension of Schwarzschild, translations along the horizon generator affine parameters single out the state known as the Hartle-Hawking or Hartle-Hawking-Israel state [4, 5], which describes physically a Hawking-radiating black hole in thermal equilibrium with a heat bath at infinity, and is the unique regular state that is invariant under all the continuous isometries [37]. A third state of interest is the Unruh state [21], defined on the union of an exterior region, the black hole region, and the black hole horizon joining the two: this state is defined in terms of positive frequency with respect to Minkowski time translations at the past infinity and with respect to the future horizon generator on the future horizon, and it describes physically a Hawking-radiating black hole with no infalling radiation from the past infinity, but it is singular on the past horizon.
On the spinless BTZ black hole spacetime, a new feature compared with Schwarzschild is that the asymptotically AdS infinity requires a boundary condition to make the theory unitary [38]. Physically, the effect of asymptotically AdS infinity on wave propagation is similar to that of a boundary at a finite distance, and the boundary condition means specifying the phase shift for waves that are reflected back from infinity. The boundary condition therefore does not allow the field to be decomposed into “incoming” and “outgoing” parts near infinity, and this prevents the construction of a time-asymmetric state similar to the Unruh state on Schwarzschild. A Boulware-type state in the BTZ black hole exterior can be readily defined, using the boundary condition at infinity and the notion of positive frequency with respect to the exterior Killing time translations, but this state is singular on the future and past horizons, like the Boulware state on exterior Schwarzschild.
The state we consider on the spinless BTZ black hole spacetime is the state induced from the $\text{AdS}_{3}$ global vacuum by the $\mathbb{Z}$ quotient described in Section III.2 [39]. We may refer to this state as a Hartle-Hawking type state: it is regular over all of the BTZ spacetime, it is invariant under the continuous isometries of the spacetime, and it describes physically a Hawking-radiating black hole in thermal equilibrium with a heat bath at asymptotic infinity. We emphasise that this state knows not just about the geometry in a single exterior region: it knows about the full global structure that the spacetime has behind the black and white hole horizons. Detectors interacting with this state have been considered in [28, 40].
On the $\mathbb{R}\text{P}^{2}$ geon and the Swedish geon, we follow the example of the BTZ spacetime and consider the state induced from the $\text{AdS}_{3}$ global vacuum by the quotients described in Sections III.3 and III.4. In each case, the state is regular on all of the spacetime, including the horizons, and the state knows about the full global spacetime structure, including that behind the horizons. What is new is that the restriction of the state to the exterior region is not invariant under the exterior Killing time translations, because the quotienting group contains elements that do not commute with these translations. Detectors interacting with this state on the $\mathbb{R}\text{P}^{2}$ geon have been considered in [12, 29, 41]. We shall consider a detector interacting with this state on the Swedish geon, and compare with detectors on the BTZ spacetime and on the $\mathbb{R}\text{P}^{2}$ geon.
IV.2 Image sum in the state induced from the $\text{AdS}_{3}$ global vacuum
@@
On each of the three quotients of an open region in $\text{AdS}_{3}$, we now consider the state induced from the $\text{AdS}_{3}$ global vacuum, as described above.
@@
We denote by $G$ the discrete group by which the quotient is taken. The Wightman function on the quotient spacetime is then given by the method-of-images sum
@@
In all three spacetimes, we consider a UDW detector on the static exterior worldline that is given in the BTZ exterior coordinates (14) by
@@ IV.3 BTZ
IV.4 $\mathbb{R}\text{P}^{2}$ geon
@@ Acknowledgements.
We thank Kirill Krasnov and Fredrik Strömberg for discussions on fundamental domains and an anonymous referee for helpful presentational suggestions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.
@@ Appendix C Topology
As the fundamental group of the punctured torus is the free group generated by the two identifications shown in Figure 8(a) [43, Section 1.2][44], it follows that $A/\{\pm 1\}$ and $B/\{\pm 1\}$ provide a pair of free generators for the action of the fundamental group on the Poincaré disc.